OpenStudy (anonymous):

Cubans is a compound composed of only c and h. It is 92.26% c by mass. What is its empirical formula

OpenStudy (anonymous):

Assume 100 g of this compound Mass C = 92.26 g mass H = 100 - 92.26=7.74 g moles C = 92.26 g/ 12.011 g/mol=7.68 moles H = 7.74 g/ 1.008 g/mol=7.68 the ratio between C and H is 1 : 1 CH = empirical formula